# Fading in the Flow: Suppression of cold gas growth in expanding galactic outflows

### v2

@@ 4.1.2 Semi-analytic multiphase wind models
Figure 13: Comparison of the FB22 multiphase wind profile (density: left, pressure: middle, velocity: right panel) due the modification to the underlying single-cloud evolution introduced in our work (in blue) with the original FB22 single-cloud model (in orange). The right panel shows different velocities, namely the wind speed $v_{r}$, the wind sound speed $c_{s}$, and the cloud speed $v_{\rm cl}$ as indicated in the legend attached to the figure. The wind parameters chosen for both the orange and blue lines are the same as that used in Fig. 6 from FB22 with an initial cloud mass of $100\ M_{\odot}$ (corresponding to $\xi_{\rm init}=3.16$ in the original FB22 model and $\xi_{\rm init}=3.95$ in our prescription). The single-cloud model (cf. Fig. 12 & Eq. E.1) that controls the mass and energy loading due to the multiphase mixing of clouds seeded in an outflow is different between the two models (in orange and blue). This highlights that the evolution of clouds in an outflow plays a crucial role in setting the large scale properties of their host outflow. These clouds become increasingly important as the efficiency of mass and energy loading in the wind increases. For reference, the gray lines show the classic single phase adiabatic wind profile (CC85). Follow the link here for a full image of all the relevant quantities similar to Fig. 6 from FB22.
@@
FB22 also developed a framework for evolving a population of cool clouds moving in a hot wind. In this framework, the single-cloud model (described in the earlier part of this section) serves as the basis for modeling the source (or sink) terms that determine the radial profile of the wind (see the equations illustrated in Fig. 3 & Eq. 17 from FB22). These source/sink terms directly affect the energy and mass loading of the outflows. Therefore, the way individual clouds interact with the background wind determines the multiphase nature of the wind they are moving in. The FB22 precription assumes a collection of cold clouds are continuously injected in the outflow. These clouds introduce energy and mass in both the cold and hot phases of the wind, depending on how they cool, mix and evaporate. To this end, we demonstrate that changes in the single-cloud evolution model can directly change the properties of the wind by solving the wind equation from FB22 (Fig. 3 & Eq. 17 in their paper) with different models of the single-cloud evolution. This is illustrated in Fig. 13, where we choose the same parameters for the single-cloud model as reported by FB22 (see Fig. 6 in their paper), shown in orange, to obtain a multiphase wind and compare it with the multiphase wind obtained when switching to our modified single-phase cloud prescription, in blue (discussed at the beginning of this section). For reference, the solution for the single phase CC85 wind is shown in gray in the figure. The wind profiles (blue and orange lines in Fig. 13) differ substantially from one another ($\gtrsim 1\ \rm dex$) over a substantial downstream distance. In order to match the simulation results, our prescription has been tuned to produce a higher mass of cold gas formed in a CC85 wind compared to the original FB22 prescription (cf. Fig. 12). Consequently, the wind in our modified prescription has more mass loading by the clouds and its properties exhibit a greater deviation from the single-phase adiabatic wind – evident in the density, pressure, and velocity profiles for the wind (blue curves showing the original FB22 prescription are closer to the gray CC85 profile). As expected, the mass in cold clouds is higher in our prescription, which results in a higher loading of mass and energy in the cold phase of the multiphase wind. This is compensated by a lower mass and energy loading into the hot phase, which is introduced into the wind as the clouds get mixed and evaporated. The higher loading of cold gas from the clouds also results in a greater reduction of entropy from the constant CC85 entropy profile beyond the region where the clouds are injected. Since cloud evolution directly affects the mass and energy loading of an outflow and its multiphase nature, winds with higher efficiency of mass loading are increasingly affected by the behavior of an increased number of clouds that are seeded in them. Overall, our comparison highlights the need to accurately test the models of individual cloud-wind evolution in order to get an accurate model of multiphase galactic outflows.
@@ 4.2 Emission from clouds in a CC85 wind
Figure 14: The temporal evolution of the total surface brightness from our cloud-crushing simulation in a CC85 wind using fiducial parameters $(\mathcal{M},\chi,t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini})=(1.496,100,0.2)$; cf. Table 1. Here, we show only a few selected snapshots that clearly demonstrate the trend that the emissivity from the cloud gradually fades with time. This is due to the weakening of radiative cooling downstream leading to a reduction in the amount of dense, cold mass growing due to shear-driven mixing in the turbulent boundary layers between the cloud and the wind. The main body of the cloud is very close to the floor temperature of our simulation, at which the emission is set to zero. This is clearly demonstrated in the inset figures showing emissivity per unit volume of the gas on a slice passing through the center of our simulation domain with normal along the $z$-axis. Emission is dominated by the turbulent boundary layers. Similar layers are integrated along the $z$-axis to obtain the projected surface brightness plots in this figure. Note that the zoom levels are different across snapshots, indicated by the scales attached along with them.
@@
Another equivalent way to interpret this is that the surface brightness of the gas depends on its emissivity, which is proportional to the square of the gas density. [^9] This is emissivity integrated across all wavelengths and might not be strictly valid for any individual emission lines like HI. Since the density of the cold gas formed due to cloud-crushing in an adiabatically expanding background keeps falling, its surface brightness is expected to be significantly diminished as the clouds move down the wind. In addition, the enthalpy flux and hence the emission can slightly differ from (the non-entrained, dense) head to (entrained, diffuse) tail, in both the homogeneous cloud-crushing and the CC85 wind because of the variation in density and $\dot{m}_{\rm cl}$ (cf. Tan et al., 2021). This is certainly not the case anymore once the cold cloud is fully entrained. Even in this state, our expanding wind setup still exhibits a strong emission gradient not present for clouds in a homogeneous background.
@@
Fig. 14 shows the surface brightness of cold gas from our fiducial cloud-crushing simulation (cf. Table 1) at different times. To calculate the surface brightness map (projected in the $z$-direction orthogonal to the wind), we calculate $n_{H}^{2}\Lambda(T)\Delta V$ for every cell in the simulation domain and add them along the $z$-direction, where $\Delta V$ is cell volume. The radiative loss for gas at the floor temperature is set to zero, i.e., $\Lambda\left(T_{\rm cl}\right)=0$. As the cloud moves in the wind, it expands, and its surface brightness diminishes, consistent with the above discussion. The insets with each of the surface brightness plots in Fig. 14 show a slice of emission rate per unit volume through the center of our simulation domain. These clearly demonstrate that emission is dominated by the radiatively cooling boundary layers between the cloud and the wind. Contributions from several such slices of emission have been integrated along the line of sight ($z$-axis, here) to generate the surface brightness maps. The bulk of the cloud is at the cooling floor of our simulation ($T=4\times 10^{4}\ \rm K$), emission from which is set to zero. Fig. 14 shows a strong gradient in surface brightness in the elongated tails of the cloud, with the surface brightness of the long cometary cloud tails dimming along the wind.
@@
We point out that the surface brightness gradient in Fig. 14 is for an individual cloud. For a more realistic comparison, a semi-analytical wind model including several clouds (cf. Section 4.1.2) has to be convolved with the global surface brightness profile. Such wind models need to be multiphase by implicitly taking into account mass and energy loading by several such clouds. Comparing predicted and observed surface brightness maps is an interesting future avenue to constrain the models and our understanding of galactic winds. In particular, a statistical measure of emission gradients can reveal the properties of the hot medium otherwise inaccessible at higher redshifts.
@@
The M82 galaxy is a classic example of a starburst-driven galaxy harboring strong outflows (Lynds & Sandage 1963), which we use as a benchmark for comparing simulation results. Clouds simulated in an expanding wind have a morphological similarity with the elongated cold gas ‘chimneys’ observed by Bolatto et al. 2024; Fisher et al. 2025 in the M82 galaxy. The infrared emission from the cold gas and dust observed using JWST in M82 by Bolatto et al. 2024; Fisher et al. 2025 is formed in a starburst-driven outflow environment, similar to what we have simulated. Long, elongated plumes of cold gas also show a similar gradient in their emission along the direction of the wind (see Figs. 4 & 6 from Fisher et al. 2025). As discussed earlier, such a strong gradient in the surface brightness of the cold clouds is not possible in ‘vanilla’ cloud-crushing because the clouds in homogeneous cloud-crushing setups lack differential expansion parallel to the wind due to the absence of a spatial gradient in the wind. It is worth noting that we are unable to reproduce the loop-like features towards the center of the wind, which are very likely magnetically dominated structures – beyond the scope of our pure hydro simulations. A detailed, quantitative comparison with the measured emission gradients can help us constrain the mass transfer rate between the phases, and in particular, study the hot, accelerated phase, which is otherwise inaccessible but crucial in order to constrain feedback models.

### v3

@@ 4.1.2 Semi-analytic multiphase wind models
Figure 14: Comparison of the FB22 multiphase wind profile (density: left, pressure: middle, velocity: right panel) due to the modification of the underlying single-cloud evolution introduced in our work (in teal) from the original FB22 single-cloud model (in yellow). The right panel shows different velocities, namely the wind speed $v_{r}$, sound speed $c_{s}$, and the cloud speed $v_{\rm cl}$ as indicated in the legend attached to the figure. The wind parameters chosen for both the yellow and teal curves are the same and mentioned in the main text. Only the single-cloud model (cf. Fig. 13 & Eq. C.1) that controls the mass and energy loading due to the multiphase mixing of clouds seeded in an outflow is different between the yellow and the teal curves. This highlights that the evolution of clouds in an outflow provides a back-reaction to the wind and plays a crucial role in setting the large scale properties of their host outflow. These clouds become increasingly important as the efficiency of mass and energy loading in the wind increases. For reference, the gray curves show the classic single phase adiabatic CC85 wind profile. Follow the link here for a full image of all the relevant quantities similar to Figs. 5 & 6 from Fielding & Bryan 2022.
Fielding & Bryan 2022 (FB22 hereafter) introduced an empirical semi-analytic model that describes the evolution of a cold cloud that starts with some initial velocity (or from rest) in a hot outflow (the ‘single-cloud’ model; cf. Section 3 in their paper). The FB22 single-cloud model provides a source/sink of energy, mass, and momentum, in a steady wind driven by supernovae in order to capture the back-reaction of clouds on their host outflow. Along with the assumption that clouds are continuously introduced at a fixed rate close to the centre of the wind, Fielding & Bryan 2022 utilised their single-cloud evolution model in developing a general framework (wind equation; cf. Fig. 3 in their paper) to predict the structure of a typical supernovae-driven multiphase galactic wind.
Although in FB22 the single-cloud model was utilised to take into account the evolution of a collection of clouds in a spatially changing background wind (that act as source terms in the wind equation), the detailed evolution of an individual cloud in a non-uniform wind, and its comparison with simulations remain largely unexplored. Since our work focuses on simulations of how individual clouds evolve in both a uniform and an expanding CC85 wind, we are able to compare the FB22 model prediction of cloud-wind interaction with our simulation results. Furthermore, the motivation is that constraining the single-cloud model using simulations in diverse wind conditions would enable the construction of realistic wind profiles. This would consequently ensure an accurate modelling of the back-reaction of the clouds in the multiphase wind equation. Prior to a comparison with FB22, it is worth noting that our simulations focus on the detailed behaviour of a cloud, considering a neighbourhood close to the cloud. The back reaction of the cloud on the volume-filling gas that comprises the wind is beyond the scope of this work. In this context, we are mostly interested in simulating the evolution of an individual cloud and do not consider the collective impact of multiple clouds that can potentially alter the large-scale properties of the wind, as discussed further.
On comparing the original FB22 single-cloud model with our simulation results, we find an unsatisfactory model prediction for the evolution of cold mass growth (see the discussion on Fig. 13 later in this section for details). [^7] For simplicity and also to compare with our simulations, we do not consider different metallicities for the cloud and the wind. Therefore, we modify the FB22 single-cloud model to better reproduce the evolution of the cold mass from our simulations. In the remainder of this section (and Appendix C), we briefly discuss the FB22 single-cloud model. Specifically, we highlight the modifications that we introduce to the original model, and motivate these modifications in order to obtain a significantly improved match with our simulation results. We also discuss tentative shortcomings of the current model and point at directions for further improvement.
At its core, the FB22 single-cloud model is a set of coupled ordinary differential equations describing the evolution of the position, velocity, and mass of a cloud in a hot wind which can be numerically solved with some initial condition (described elaborately in Appendix C). One of the most important components in this model is radiative cooling of the turbulent mixing layer which forms as the cloud moves and mixes in the wind. This essentially captures the growth/destruction of cold clouds in moving in a hot wind. Fig. 13 shows a comparison of our numerical simulations – both in an expanding CC85 (teal curves) as well as uniform plane-parallel wind (yellow curves) – with the prediction of the FB22 single-cloud model (shown in thick, solid lines with a lighter shade in Fig. 13). One can see that the numerical results do not match the model sufficiently, with $\gtrsim 1$ dex more cold gas mass in the simulations compared to the model. In addition, the FB22 model is also unable to reproduce the suppression of cold mass growth of clouds in an expanding wind when compared to a uniform plane-parallel wind – a definitive trend in our simulation. This is concerning, since the FB22 is used as a basis for multiphase sub-grid model of galactic winds with the aim of being eventually incorporated into cosmological simulations (Smith et al. 2024a, b) as well as to interpret observations (Lopez et al., 2025; Xu et al., 2023). This mismatch in the cold gas mass also propagates to the overall wind properties – as we discuss below.
As a cloud moves down the wind, it can be either in the slow ($\xi\leq 1$) or in the fast ($\xi>1$) cooling regime (cf. Eq. C.1), depending on a competition between radiative cooling and turbulent mixing. Simulations have found that the growth rate of the clouds are different in these two regimes (Ji et al. 2019; Tan et al. 2021) and is interpreted as fast cooling causing a change in the fractal nature of the turbulent boundary layer around cold clouds (Sreenivasan et al. 1989; Fielding et al. 2020). An important point to note is that the difference between the original model prediction and the simulation results increase with time as clouds become increasingly co-moving with the wind. This might be because the FB22 single-cloud model only considers shear-driven hydrodynamic instabilities (Mandelbrot 1975; Federrath et al. 2009; Fielding et al. 2020; Tan & Fielding 2024) to be causing the turbulent mixing. Over time, as shear diminishes, the model predicts lower production of efficiently cooling mixed gas, resulting in a reduced growth of cold gas in a cloud. However, simulations have routinely shown that cold clouds always maintain a turbulent boundary layer and keep growing in mass even when they are co-moving with the background wind. Once entrained, boundary layers around cold clouds can be maintained by pulsations (Waters & Proga 2019; Gronke & Oh 2023), or local cooling flows (Tan & Oh 2021; Dutta et al. 2022).
To support this speculation, we show in Fig. 13, using solid dark coloured curves, a new model prediction which produces a substantially improved match with the simulation data. Motivated from the semi-analytic work by Nikolis & Gronke 2024 (see also Tan et al., 2023), we propose that the inflow velocity due to cooling, when shear diminishes depends, does not diminish to zero but instead depends on the sound speed in the cold phase. This modification to the model makes it reproduce the general trend of mass growth in the simulations, in particular the suppression of cold mass in clouds in an expanding CC85 wind compared to a uniform one. It is worth noting that this is an empirical prescription and further work is needed to build an adequate understanding of the underlying physics that maintain a turbulent layer around cold clouds after entrainment.
The exact modification to the original FB22 single-cloud model is as follows (see Eq. C.1 for details). Instead of defining the inflow velocity $v_{\rm in}$ that causes mass growth in the cloud as $v_{\rm in}=f_{\rm turb}|v_{\rm rel}|\xi^{\alpha}$, where $v_{\rm rel}=v_{\rm wind}-v_{\rm cl}$, we re-define inflow velocity
$v_{\rm in}={\rm max}\left[f_{\rm turb}v_{\rm rel}\xi^{\alpha},f_{\rm sat}c_{\rm s,cl}\left(\frac{l^{\prime}}{c_{\rm s,cl}t_{\rm cool,cl}}\right)^{1/4}\right].$ (4.1)
We found that in the case of a uniform wind (‘vanilla’), $f_{\rm sat}=0.075$ and $l^{\prime}$ set to the initial cloud size best reproduces the simulation results (which is in agreement with previous simulations; see Gronke & Oh, 2020; Fielding et al., 2020; Tan et al., 2021 and references therein for details on the $1/4$ scaling). However, in case of a CC85 wind, $f_{\rm sat}=0.19$ and $l^{\prime}=R_{\rm cl}=(3M_{\rm cl}/4\pi\rho_{\rm cl})^{1/3}$ gives a good match with simulations. Since $l^{\prime}$ represents the characteristic length scale for turbulence in the boundary layers, an expression for $l^{\prime}$ that scales with cloud size is intuitively expected. Nevertheless, the fact that we need different values for the free parameters for different wind conditions reveals the limitation of the current model to accurately capture the cloud geometry and its evolution in diverse cloud-wind conditions.
One possibly limiting assumption in the model is that a cloud instantaneous stretches along the wind as it cools and it is determined by a constant factor $f_{\rm cool}$ in the FB22 model (see Appendix C for details). Clouds can however undergo differential expansion and stretch over time in directions both along and orthogonal to the wind (Huang et al. 2020; Nikolis & Gronke 2024), as also seen in our simulations. The model can, therefore, be improved by position/time dependent free parameters to better capture the cloud geometry and its evolution downstream. The free parameters listed here are, therefore, not robust across diverse wind or cloud properties (and would need to be adjusted depending on the wind+cloud parameters) in accurately predicting the mass evolution nor a representative best fit value. Instead, the values for these free parameters used here is intended as a demonstration in support of the idea that there is a need to floor the inflow velocity (that causes the mass growth in cold clouds) and not to let it diminish over time. Note that the agreement between the original FB22 model and our modified prescription is significantly better in the cloud velocity than in mass. In fact, the cloud velocity in our ‘modified’ model matches the original FB22 model on a percent level. We detail this further in Appendix C. A detailed exploration of the model’s parameter space as well as developing a more robust physical accurate model is beyond the scope of this work and is left for future studies. Our primary goal is to emphasize the need for improving existing models by calibrating them against cloud-crushing simulations in evolving background wind conditions. In the remainder of this section, we discuss that changes in the single-cloud model can potentially result in non-negligible differences in the multiphase nature of their host outflows.
@@
Fielding & Bryan 2022 also developed a framework for evolving a population of cool clouds moving in a hot wind. In this framework, the single-cloud model (described in the earlier part of this section) serves as the basis for modelling the source (or sink) terms that determine the radial profile of the wind (see the equations illustrated in Fig. 3 & Eq. 17 from FB22). These source/sink terms directly affect the energy and mass loading of the outflows. Therefore, the way individual clouds interact with the background wind gives rise to a back reaction on their host outflow and determines the multiphase nature of the wind. The FB22 prescription assumes a collection of cold clouds that are continuously injected in the outflow. These clouds introduce energy and mass in both the cold and hot phases of the wind, depending on how they cool, mix and evaporate. To this end, we demonstrate that changes in the single-cloud evolution model can directly change the properties of the wind by solving the wind equation from FB22 (Fig. 3 & Eq. 17 in their paper) with different models of the single-cloud evolution. This is illustrated in Fig. 14, where we compare the FB22 single-cloud model, set as source/sink terms in the wind equation, with our single-cloud model. Specifically, we solve the FB22 wind equation with $5\ \rm M_{\odot}yr^{-1}$ star formation and energy and mass loading efficiencies of $0.5$ and $0.2$ respectively. The coupling efficiency of the cloud with the wind $\eta_{\rm M,cold}$ is set to $0.15$. We set the initial cloud mass to $100\ \rm M_{\odot}$ ($\xi_{\rm ini}=2.63$) introduced at about $220\ \rm pc$ with a temperature of $8\times 10^{3}\ \rm K$ and an initial radial velocity of $30\ \rm km\ s^{-1}$ in a steady CC85 wind which has a sonic radius of $200\ \rm pc$ opening over a full $4\pi$ steradians in solid angle. Additionally, the gravity is set by a circular velocity of $150\ \rm km\ s^{-1}$. The wind and the cloud metallicities are set to $2\ Z_{\odot}$ and $Z_{\odot}$, respectively. In the Fig. 14, we compare the wind profiles on changing from the original FB22 single-cloud model (in yellow) to our ‘modified’ single-cloud model (in teal). The two wind profiles differ substantially from one another ($\sim 1\ \rm dex$ in pressure and $\sim 0.5\ \rm dex$ in density and velocity) over almost their entire downstream distance ($\gtrsim 1\rm\ kpc$).
Since cloud evolution directly affects the mass and energy loading of an outflow and its multiphase nature, winds with higher efficiency of mass loading are increasingly affected by the behaviour of an increased number of clouds that are seeded in them. It is worth noting that our discussion in this section is far from being a comprehensive modelling of multiphase galactic outflows. We simply demonstrate that wind properties can significantly influence the evolution of embedded clouds, which in turn modify the large-scale structure of the host outflow – setting up a self-adjusting interplay that further shapes subsequent cloud evolution. Overall, our comparison highlights the need to accurately test models of individual cloud-wind evolution in order to get an accurate model of multiphase galactic outflows. Such accurate models can greatly improve sub-grid prescriptions of cold gas in cosmological simulations or simulations with limited resolutions (Weinberger & Hernquist 2023; Smith et al. 2024a; Butsky et al. 2024).
@@ 4.2 Emission from clouds in a CC85 wind
Figure 15: The temporal evolution of the total surface brightness from our cloud-crushing simulation in a CC85 wind using fiducial parameters $(\mathcal{M},\chi,t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini})=(1.496,100,0.2)$; cf. Table 1. Here, we show only a few selected snapshots that clearly demonstrate the trend that the emissivity from the cloud gradually fades with time. This is due to the weakening of radiative cooling downstream leading to a reduction in the amount of dense, cold mass growing due to shear-driven mixing in the turbulent boundary layers between the cloud and the wind. The main body of the cloud is very close to the floor temperature of our simulation, at which the emission is set to zero. This is clearly demonstrated in the inset figures showing emissivity per unit volume of the gas on a slice passing through the centre of our simulation domain with normal along the $z$-axis. Emission is dominated by the turbulent boundary layers. Similar layers are integrated along the $z$-axis to obtain the projected surface brightness plots in this figure. Note that the zoom levels are different across snapshots, indicated by the scales attached along with them.
@@
Another equivalent way to interpret this is that the surface brightness of the gas depends on its emissivity, which is proportional to the square of the gas density. [^8] This is emissivity integrated across all wavelengths and might not be strictly valid for any individual emission lines like HI. Since the density of the cold gas formed due to cloud-crushing in an adiabatically expanding background keeps falling, its surface brightness is expected to be significantly diminished as the clouds move down the wind. In addition, the enthalpy flux and hence the emission can slightly differ from (the non-entrained, dense) head to (entrained, diffuse) tail, in both the homogeneous cloud-crushing and the CC85 wind because of the variation in density and $\dot{m}_{\rm cl}$ (cf. Tan et al., 2021). This is certainly not the case once the cold cloud is fully entrained. Even in this state, our expanding wind setup still exhibits a strong emission gradient not present for clouds in a homogeneous background.
@@
Fig. 15 shows the surface brightness of cold gas from our fiducial cloud-crushing simulation (cf. Table 1) at different times. To calculate the surface brightness map (projected in the $z$-direction orthogonal to the wind), we calculate $n_{H}^{2}\Lambda(T)\Delta V$ for every cell in the simulation domain and add them along the $z$-direction, where $\Delta V$ is cell volume. The radiative loss for gas at the floor temperature is set to zero, i.e., $\Lambda\left(T_{\rm cl}\right)=0$. As the cloud moves in the wind, it expands, and its surface brightness diminishes, consistent with the above discussion. The insets with each of the surface brightness plots in Fig. 15 show a slice of emission rate per unit volume through the centre of our simulation domain. These clearly demonstrate that emission is dominated by the radiatively cooling boundary layers between the cloud and the wind. Contributions from several such slices of emission have been integrated along the line of sight ($z$-axis, here) to generate the surface brightness maps. The bulk of the cloud is at the cooling floor of our simulation ($T=4\times 10^{4}\ \rm K$), emission from which is set to zero. Fig. 15 shows a strong gradient in surface brightness in the elongated tails of the cloud, with the surface brightness of the long cometary cloud tails dimming along the wind.
@@
We point out that the surface brightness gradient in Fig. 15 is for an individual cloud. For a more realistic comparison, a semi-analytical wind model including several clouds (cf. Section 4.1.2) has to be convolved with the global surface brightness profile. Such wind models need to be multiphase by implicitly taking into account mass and energy loading by several such clouds. Comparing predicted and observed surface brightness maps is an interesting future avenue to constrain the models and our understanding of galactic winds. In particular, a statistical measure of emission gradients can reveal the properties of the hot medium otherwise inaccessible at higher redshift.
@@
The M82 galaxy is a classic example of a starburst-driven galaxy harbouring strong outflows (Lynds & Sandage 1963), which we use as a benchmark for comparing simulation results. Clouds simulated in an expanding wind have a morphological similarity with the elongated cold gas ‘chimneys’ observed by Bolatto et al. 2024; Fisher et al. 2025 in the M82 galaxy. The infrared emission from the cold gas and dust observed using JWST in M82 by Bolatto et al. 2024; Fisher et al. 2025 is formed in a starburst-driven outflow environment, similar to what we have simulated. Long, elongated plumes of cold gas also show a similar gradient in their emission along the direction of the wind (see Figs. 4 & 6 from Fisher et al. 2025). As discussed earlier, such a strong gradient in the surface brightness of the cold clouds is not possible in ‘vanilla’ cloud-crushing because the clouds in homogeneous cloud-crushing setups lack differential expansion parallel to the wind due to the absence of a spatial gradient in the wind. It is worth noting that we are unable to reproduce the loop-like features towards the centre of the wind, which are very likely magnetically dominated structures – beyond the scope of our pure hydro simulations. A detailed, quantitative comparison with the measured emission gradients can help us constrain the mass transfer rate between the phases, and in particular, study the hot, accelerated phase, which is otherwise inaccessible but crucial in order to constrain feedback models.